### v2

# Central Dogma Transformer II An AI Microscope for Understanding Cellular Regulatory Mechanisms

Abstract
[…]
Introduction
The central dogma1 describes cellular information flow across three molecular layers: DNA encodes genetic instructions, RNA transmits and regulates this information, and proteins execute cellular functions. Artificial intelligence has transformed our ability to model each layer individually2–7, yet a fundamental limitation remains: current models lack interpretability8,9. Their internal representations do not correspond to biological entities or relationships that researchers can examine and validate. The goal of biological research, however, is not merely to predict cellular responses but to understand the mechanisms that produce them. What biology needs is not task-oriented AI optimizing predictions, but mechanism-oriented AI revealing regulatory structure.
To address this limitation, we reframe biological AI as an “AI microscope” whose attention maps are directly interpretable as regulatory structure. The Central Dogma Transformer10 implements this principle by mirroring the central dogma in its architecture: DNA self-attention captures genomic relationships, RNA self-attention reflects gene co-regulation, and DNA-to-RNA cross-attention models transcriptional control. Because each attention mechanism corresponds to a specific biological relationship, the resulting maps provide direct readouts of regulatory organization. This architectural inductive bias fundamentally changes the meaning of prediction accuracy: unlike models optimizing prediction as an end goal, CDT-II uses prediction accuracy as evidence that the model has learned correct regulatory structure. As training progresses, improving predictions and discovering biology become the same objective.
Critically, CDT-II’s success stems not only from model architecture but also from task formulation. The model receives raw per-cell expression values as input and predicts expression changes (log2 fold changes) as output—without being provided the difference or any reference to the unperturbed state. This design forces the model to learn what constitutes “change” and which genes influence which: to predict how gene B changes when locus A is perturbed, the model must internalize the regulatory relationship between A and B. By not explicitly computing differences in the input, we ensure that attention maps reflect genuinely learned regulatory structure rather than artifacts of input preprocessing.
While CDT v1 required three pre-trained language models, CDT-II reduces the input requirements to genomic sequence embeddings and raw per-cell expression, enabling experimental biologists to observe regulatory networks in their own data. CDT-II is designed as an integration platform rather than a standalone model: pre-trained language models provide the best available representations of each molecular layer, while task-driven training on experimental data organizes this knowledge into interpretable regulatory maps. The DNA embedding component is deliberately modular, allowing it to be upgraded as genomic foundation models improve. Because each attention mechanism corresponds to a specific biological relationship, CDT-II’s outputs generate experimentally testable hypotheses—creating a feedback cycle in which experimental validation improves the model, and the improved model generates sharper hypotheses for the next round of experiments.
We validated CDT-II on a large-scale K562 CRISPRi screen11,12, holding out five perturbation targets entirely—including GFI1B, a master transcriptional regulator with hundreds of experimentally characterized trans-regulatory targets12—to test whether regulatory networks could be recovered from unsupervised attention patterns. Here we report five principal findings. First, CDT-II achieves a validation Pearson correlation of 0.64 (per-gene mean $r=0.84$) without requiring RNA or protein language models. Second, attention analysis recovers GFI1B’s trans-regulatory network with 6.6-fold enrichment over random expectation ($P=3.5\times 10^{-17}$), demonstrating unsupervised discovery of a known regulatory program. Third, two distinct attention mechanisms—RNA self-attention and DNA-to-RNA cross-attention—independently identify an overlapping RNA processing module (80% gene overlap) with highly significant functional enrichment (RNA binding, $P=1\times 10^{-16}$), revealing hierarchical regulation across central dogma layers. Fourth, an ablation study establishes that gene set quality, determined by cross-dataset reproducibility, governs model resolution. Fifth, systematic comparison against ENCODE K562 regulatory annotations reveals that cross-attention autonomously focuses on known regulatory elements, including DNase hypersensitive sites and CTCF binding sites, across all five held-out genes. Together, these results demonstrate that CDT-II’s attention maps provide biologically meaningful observations of regulatory networks.
Results
CDT-II architecture and the AI microscope framework
CDT-II implements a two-modality architecture that follows the directional logic of the central dogma (Fig. 1). Genomic DNA embeddings, generated by Enformer3 from sequences centered on perturbation sites, are represented as a $896\times 3{,}072$ matrix (896 genomic bins spanning approximately 115 kilobases at 128-base-pair resolution, each with a 3,072-dimensional embedding). Per-cell RNA expression values for 2,361 genes are encoded by a RawExpressionEncoder that combines learned gene identity embeddings with projected expression values, producing gene-level representations of dimension 512.
Each modality first passes through self-attention layers13 that capture internal structure: two layers for DNA (learning relationships between genomic positions) and one layer for RNA (learning co-regulatory relationships between genes). A cross-attention layer then models transcriptional regulation, with RNA representations serving as queries and DNA representations as keys and values—enabling the model to learn which genomic regions are relevant for each gene’s regulation. All attention operations use 8 heads and preserve the original dimensionality, ensuring that attention maps remain directly interpretable as gene-by-gene or gene-by-position matrices.
After cross-attention, a Virtual Cell Embedding (VCE)10 module integrates DNA and RNA representations into a unified 512-dimensional cell-state vector. This integration uses learned attention pooling, which weights each modality’s elements based on their relevance. A two-layer task head then projects this VCE to predict log2 fold changes for all 2,361 genes simultaneously. The complete model contains approximately 21 million parameters (see Online Methods for training details).
The dataset comprised 15,657 K562 cells from the CRISPRi screen12: 8,250 cells from TSS perturbations (27 genes) and 7,407 cells from SNP perturbations (420 loci). Of these, 13,620 cells were used for training and 2,037 cells were reserved for validation. Training included 6,213 TSS cells (22 genes) and all 7,407 SNP cells. Validation comprised five genes held out entirely at the gene level: GFI1B, CD52, TFRC, CD44, and TNFSF9. This gene-level holdout ensures that validation performance reflects generalization to entirely unseen perturbation targets—a stringent test of the model’s ability to learn transferable regulatory patterns rather than memorize gene-specific effects.
The AI microscope framework assigns distinct roles to each component: prediction accuracy is the focus adjustment, the gene set is the objective lens, and attention maps are the image. This framing structures the subsequent analyses: we first validate the lens quality through ablation (Fig. 2), confirm proper focus through prediction performance (Fig. 3), and then examine what the microscope reveals (Fig. 4–6).
Figure 1: CDT-II architecture and interpretable attention maps. Unlike conventional deep learning models whose internal representations are opaque, CDT-II produces attention maps that directly correspond to biological relationships. CDT-II processes two input modalities following the central dogma. Left branch: Genomic DNA sequences centered on perturbation loci ($\pm$57 kb) are encoded by Enformer, projected to a common dimension, and processed by two DNA self-attention layers that capture genomic relationships between positions. Right branch: Per-cell RNA expression values from scRNA-seq data are encoded by the RawExpressionEncoder, projected, and processed by one RNA self-attention layer that captures gene co-regulatory relationships. Center: A DNA-to-RNA cross-attention layer models transcriptional control, with RNA representations as queries and DNA representations as keys and values. The Virtual Cell Embedder integrates both modalities through attention pooling into a unified cell-state vector, which the Task Head projects to predict perturbation effects for all 2,361 genes. Right panels: Each attention mechanism produces an interpretable map—DNA self-attention reveals genomic relationships (blue), RNA self-attention reveals gene co-regulation networks (orange), and cross-attention reveals transcriptional control patterns (purple). These maps constitute the primary output of the “AI microscope,” enabling direct observation of regulatory structure.
Gene set quality determines microscope resolution
To justify the choice of 2,361 genes—and to establish a general principle for applying CDT-II to new datasets—we conducted an ablation study across gene set configurations (Fig. 2).
Our initial attempt used 9,335 genes detected in at least 10% of cells in the CRISPRi dataset12, applying only a single-dataset expression threshold. This configuration ($\sim$54 million parameters) plateaued at a validation Pearson $r$ of 0.37, and attention maps exhibited collapse—a diffuse, uninformative pattern indicating that the model could not resolve regulatory structure from the noisy gene set.
Based on this result, we hypothesized that genes lacking reproducibility across independent experiments introduced noise that degraded model performance. We therefore applied a cross-dataset filter, retaining only the 2,360 genes detected as expressed in two independent K562 CRISPRi screens12,14, and added GFI1B as a critical test case (2,361 total, $\sim$21 million parameters). The effect was immediate: with reduced noise, the learning curve improved steadily throughout training, reaching $r=0.64$ with training r=0.65r=0.65—indicating minimal overfitting.
The contrast is stark: nearly quadrupling the gene count degraded performance by 42%, while filtering to cross-dataset reproducible genes yielded a 73% improvement. Notably, the larger model ($\sim$54 million parameters) underperformed the smaller one ($\sim$21 million parameters), the opposite of what would be expected if model capacity were the limiting factor. This establishes a counterintuitive but practically important principle: gene set quality, not quantity, determines the resolution of the AI microscope. Cross-dataset reproducibility filtering removes genes whose perturbation responses are dominated by technical or biological noise, enabling the model to learn sharper regulatory relationships.
[…]
CDT-II predicts cell-level perturbation effects
Having established the gene set, we evaluated CDT-II’s predictive performance (Fig. 3).
CDT-II achieved an overall validation Pearson $r$ of 0.64 ($R^{2}=0.41$) and Spearman $\rho$ of 0.46, predicting cell-level log2 fold changes across all five held-out genes simultaneously.
To evaluate how well CDT-II captures the trans-regulatory “fingerprint” of each perturbation, we averaged predictions and observations across cells for each held-out gene and computed Pearson $r$ on these mean trans-effect profiles: GFI1B ($r=0.88$), TNFSF9 ($r=0.86$), TFRC ($r=0.85$), CD44 ($r=0.85$), and CD52 ($r=0.75$), yielding a mean $r$ of 0.84. This high pseudo-bulk correlation demonstrates that CDT-II faithfully captures the transcriptome-wide pattern of perturbation effects—the regulatory “fingerprint” that distinguishes one perturbation from another.
To contextualize these correlations, we examined the empirical reproducibility of CRISPRi perturbation effects across independent studies. Morris et al. found that only 56% of GFI1B trans-targets identified in their STING-seq screen replicated in the independent Gasperini dataset12,14, and direct comparison of pseudo-bulk effect sizes between the two K562 CRISPRi studies yields $r\approx 0.79$. CDT-II’s pseudo-bulk correlation (mean $r=0.84$) surpasses this empirical ceiling, indicating that the model captures essentially all reproducible biological signal. The lower cell-level correlation ($r=0.64$) reflects stochastic variation inherent to single-cell measurements—noise that is averaged out at the pseudo-bulk level but represents a fundamental limit of single-cell prediction.
Trans-target prediction provided a stringent test of CDT-II’s ability to capture regulatory wiring. Among the top 50 predicted trans-targets, 26 overlapped with the top 50 experimentally determined GFI1B targets (52%, approximately 25-fold enrichment over the 2.1% expected by chance), and CD52 showed comparable overlap (25 of 50, 50%). Notable GFI1B trans-targets among these include G0S2, TXNIP, and JUN—genes involved in cell cycle arrest, oxidative stress response, and transcriptional regulation, respectively, consistent with GFI1B’s known functions in hematopoietic differentiation.
A
B
C
Figure 3: CDT-II predicts cell-level perturbation effects. (A) Overall scatter plot of predicted versus actual log2 fold changes across all validation cells (n=2,037n=2{,}037 cells $\times$ 2,361 genes), showing Pearson $r=0.64$. Each point represents a cell-gene pair, with density indicated by color intensity. (B) Per-gene prediction performance for five validation genes (GFI1B, TNFSF9, TFRC, CD44, CD52), with pseudo-bulk correlations (mean across cells) ranging from 0.75 to 0.88 (mean $r=0.84$). (C) Trans-effect profile for GFI1B: mean predicted versus mean actual effects across all 2,361 genes, showing strong recovery of the perturbation-specific regulatory signature ($r=0.88$).
Attention maps reveal the GFI1B regulatory network
We next examined what CDT-II’s attention maps reveal about GFI1B—a transcription factor held out entirely from training (Fig. 4).
The attention analysis follows the central dogma flow. DNA self-attention (Fig. 4A) reveals how the model learns relationships between genomic positions within the 115 kb window centered on the perturbation site. DNA-to-RNA cross-attention (Fig. 4B) shows how each gene attends to specific genomic regions, capturing transcriptional regulatory relationships. RNA self-attention (Fig. 4C–D) captures co-regulatory relationships between genes, revealing the gene regulatory network structure.
The RNA self-attention matrix $2{,}361\times 2{,}361$] captures learned co-regulatory relationships between genes. Extracting the GFI1B query row—representing which genes GFI1B “attends to” during inference—revealed a biologically coherent ranking. Cell cycle regulators including CDCA8, CDC20, KIF2C, and KIF14 appeared in the top ranks, consistent with GFI1B’s established role in cell cycle control during hematopoietic differentiation15.
Visualizing the top 5% of attention edges as a network (Fig. 5A) reveals GFI1B as a hub gene with multiple connection types: downstream targets (genes that GFI1B regulates), upstream regulators (genes that influence GFI1B), and bidirectional relationships suggesting reciprocal regulatory dynamics. This bidirectional structure captures the feedback loops characteristic of transcription factor networks.
To quantify the correspondence between attention-derived and experimentally determined regulatory networks, we ranked all 2,361 genes by GFI1B attention weight and, independently, by absolute experimental effect size (mean $|\log_{2}\text{FC}|$ upon GFI1B perturbation12), then measured overlap between the top $N$ genes from each ranking. At top 100, 28 genes appeared in both lists, representing a 6.6-fold enrichment over random expectation ($P=3.5\times 10^{-17}$ by hypergeometric test). This demonstrates that CDT-II’s attention maps recover known regulatory relationships without supervision.
Cell-to-cell variation in attention patterns revealed state-dependent regulatory networks (Fig. 4E). Across validation cells, pairwise correlations of GFI1B attention vectors showed a mean of $r=0.75$, indicating that while the core network structure is preserved, CDT-II modulates attention weights based on each cell’s expression state. This per-cell variation is a direct consequence of using raw expression as input—each cell’s unique transcriptional profile generates a unique view through the AI microscope.
A
B
Figure 4: Attention maps reveal regulatory structure. (A) DNA self-attention patterns across the 115 kb genomic window for two layers. (B) DNA-to-RNA cross-attention profiles for five TSS perturbations.
C
D
E
Figure 4 (continued). (C) RNA self-attention per-head analysis for GFI1B, showing that different attention heads capture distinct aspects of the regulatory network. (D) RNA self-attention 2D heatmap showing gene-gene co-regulatory relationships centered on GFI1B and its top-attended genes. (E) Cell-to-cell attention variation. Left: pairwise correlations of GFI1B attention vectors for 20 representative cells sampled from 477 validation cells (mean $r=0.75$ across all pairs), showing that core network structure is preserved while attention weights are modulated by each cell’s expression state. Right: genes with the most variable attention across cells, representing cell-state-dependent regulatory targets.
Convergent attention reveals an RNA processing module
A notable finding emerged from comparing regulatory modules detected by CDT-II’s two distinct attention mechanisms.
We applied Louvain community detection16 to networks derived from each mechanism separately. From the RNA self-attention matrix, we identified community C1 (679 genes) with highly significant enrichment for RNA binding (Gene Ontology17 molecular function; $P=5\times 10^{-16}$) and metabolism of RNA (Reactome18; $P=2\times 10^{-9}$). From the DNA-to-RNA cross-attention matrix—where gene similarity was computed based on shared attention to the same genomic regions—we identified community C3 (165 genes) with even stronger enrichment for RNA binding ($P=1\times 10^{-16}$), spliceosome (KEGG19; $P=4\times 10^{-10}$), and mRNA splicing (Reactome18; $P=2\times 10^{-9}$).
The convergence between these independently derived modules was striking: 132 of 165 cross-attention genes (80%) were contained within the 679-gene self-attention community (2.8-fold enrichment over the 47.5 expected by chance; hypergeometric $P=9.3\times 10^{-46}$). This core module of 132 genes spans multiple levels of RNA processing: spliceosome components (SF3A3, SF3B4, PRPF3, SNRNP27, SNRPG, BCAS2, RBM8A), RNA modification and decay factors (YTHDF2 for m6A reading, ADAR for RNA editing, EXOSC10 for exosome-mediated degradation, TUT4 for uridylation, XPO1 for nuclear export), chromatin regulators (HDAC1, KDM1A/LSD1, ARID1A of the SWI/SNF complex), transcription machinery (CDC73 of the PAF complex, TAF12 of TFIID, TRIM33), and translation initiation factors (EIF2B3, EIF2B4, EIF4G3).
The biological interpretation of this convergence is hierarchical. RNA self-attention captures post-transcriptional co-regulation—genes whose expression levels co-vary in response to perturbation, reflecting shared regulatory mechanisms at the mRNA level. DNA-to-RNA cross-attention captures transcriptional co-regulation—genes that attend to the same genomic regions, suggesting shared upstream regulators encoded in DNA.
That both mechanisms independently identify the same RNA processing module indicates that these genes are coordinately regulated at two distinct levels of the central dogma: they share transcriptional regulatory inputs (detected by cross-attention) and participate in the same post-transcriptional feedback networks (detected by self-attention) (Fig. 5B–C). Notably, while the architecture provides separate attention mechanisms, this convergence was not explicitly encouraged—it emerged from training on perturbation data alone.
A
B
C
Figure 5: GFI1B regulatory network and convergent attention on an RNA processing module. (A) Network visualization of GFI1B regulatory relationships derived from the top 5% of RNA self-attention edges. GFI1B (red, center) functions as a hub gene connected to downstream targets (blue), upstream regulators (orange), and genes with bidirectional relationships (purple). Edge thickness is proportional to attention weight. (B) GO enrichment comparison showing both attention mechanisms identify “RNA Binding” as the top enriched term. RNA self-attention community C1 (679 genes) and DNA-to-RNA cross-attention community C3 (165 genes) show highly significant enrichment for RNA processing functions. (C) Venn diagram showing 132-gene core overlap (80% of cross-attention community). Core genes span spliceosome components, RNA processing factors, chromatin regulators, and transcription machinery.
Cross-attention recovers known regulatory elements
To test whether CDT-II’s cross-attention learns genuine regulatory structure, we compared attention patterns against ENCODE regulatory annotations20 for K562 cells (Fig. 6).
Enformer3, which provides CDT-II’s DNA embeddings, was pre-trained on epigenomic tracks including DNase-seq and histone ChIP-seq, so the initial embeddings already encode regulatory element positions. However, cross-attention does not operate on these raw embeddings: a projection layer and two DNA self-attention layers, all trained end-to-end on perturbation prediction, substantially transform the representations before cross-attention computes gene-specific regulatory maps (see Discussion).
For each of the five held-out genes, we extracted the mean cross-attention profile across validation cells, identifying the top 10% of attended genomic bins (89 of 896 bins spanning the 115 kb window). We then overlapped these high-attention bins with ENCODE K562 peak calls for five regulatory marks: DNase I hypersensitivity (open chromatin), CTCF binding, H3K27ac (active enhancers), H3K4me1 (poised enhancers), and H3K4me3 (active promoters). Fisher’s exact tests with Haldane correction were applied to each gene–mark combination (24 testable of 25; one combination, CD44 $\times$ H3K27ac, yielded no ENCODE peaks in the window).
[…]
To assess statistical robustness, we performed circular permutation tests (n=1,000n=1{,}000), shifting attention profiles along the genomic axis while preserving their autocorrelation structure. Among the 15 combinations tested (DNase, CTCF, and H3K4me3 $\times$ 5 genes), 14 achieved permutation $P<0.005$ (Supplementary Fig. 1a), confirming that the observed enrichments are not artifacts of attention smoothness or genomic feature clustering. Threshold robustness analysis showed that enrichments remained significant across a range of attention thresholds (top 5–20%), with effect sizes varying smoothly rather than appearing only at a single arbitrary cutoff (Supplementary Fig. 1b).
To further characterize what cross-attention captures, we classified each genomic bin by its ENCODE annotation—promoter (H3K4me3+), active enhancer (H3K27ac+ & H3K4me3-), CTCF-only, or unannotated—and compared attention weights across classes (Fig. 6B). Bin class separation was substantial: CTCF bins received significantly higher attention than unannotated bins (Cohen’s $d=2.4$), as did promoter bins ($d=1.4$). Kruskal–Wallis tests confirmed highly significant differences across all bin classes ($P=1\times 10^{-45}$). Notably, CTCF-associated bins showed the most consistent gene-specificity signal, ranking as the most informative mark across all five genes by entropy analysis.
These results demonstrate that CDT-II’s task-driven training selects and organizes genomic information into gene-specific regulatory maps. The cross-attention mechanism, optimized solely to predict perturbation effects, converges on the same regulatory grammar catalogued by ENCODE—providing strong evidence that the AI microscope reveals genuine regulatory biology rather than statistical artifacts.
Figure 6: Cross-attention recovers known regulatory elements. (A) Enrichment heatmap showing odds ratios (Fisher’s exact test) for overlap between CDT-II’s top 10% cross-attention bins and ENCODE K562 regulatory peaks across five held-out genes and five marks. 23 of 24 testable combinations achieve $P<0.001$ (CD44 $\times$ H3K27ac yielded no peaks). (B) Violin plots of attention weights stratified by ENCODE bin class (promoter, active enhancer, CTCF, unannotated); black lines indicate means, red lines indicate medians. CTCF-associated bins show the largest separation from unannotated bins (Cohen’s $d=2.4$), with Kruskal–Wallis $P=1\times 10^{-45}$ across all classes. See Supplementary Fig. 1 for circular permutation tests and threshold robustness analysis.
Discussion
[…]
We note that Enformer’s pre-training on epigenomic tracks across many cell types means the initial DNA embeddings encode general regulatory element positions based on sequence alone—Enformer receives no cell-type information and produces identical embeddings regardless of cellular context. These cell-type-agnostic embeddings are then substantially transformed before cross-attention—through a learned projection layer and two DNA self-attention layers, all trained on K562 perturbation prediction—so the cross-attention maps reflect CDT-II’s cell-type-specific, task-specific understanding rather than Enformer’s general-purpose representations. The key finding is that this multi-layer transformation selects the gene-specific regulatory elements relevant to each perturbation target in K562, transforming Enformer’s cell-type-agnostic genomic representation into an interpretable, gene-resolved regulatory map that aligns with K562 ENCODE annotations20. More broadly, CDT-II is designed as an integration platform: its goal is not to outperform any individual component, but to combine pre-trained representations with experimental perturbation data and extract interpretable regulatory structure that neither source can provide alone. The DNA embedding module is deliberately modular—Enformer can be replaced with newer models such as AlphaGenome21 or Evo22—so the relevant question is not how much information originates from any single component, but whether the integrated system produces biologically meaningful outputs.
The prominence of CTCF in cross-attention maps has implications for understanding three-dimensional chromatin organization. CTCF is a primary architectural protein that establishes topologically associating domain (TAD) boundaries and chromatin loop anchors, compartmentalizing the genome into regulatory neighborhoods23. The finding that CTCF-associated bins receive the highest attention weights across all five held-out genes (Cohen’s $d=2.4$) and show the most consistent gene-specificity signal suggests that CDT-II may be learning aspects of three-dimensional regulatory structure from one-dimensional sequence information alone. This is particularly striking because the model receives no explicit chromatin conformation data—neither Hi-C contact maps nor TAD annotations. Future integration of chromatin conformation data could further sharpen this aspect of the AI microscope’s resolution, potentially enabling direct visualization of gene-specific regulatory neighborhoods.
The ablation study establishes a practically important principle: gene set quality, determined by cross-dataset reproducibility filtering, governs the resolution of the AI microscope. Training with 9,335 genes produced inferior predictions ($r=0.37$), while 2,361 cross-dataset validated genes yielded both superior accuracy ($r=0.64$) and interpretable attention maps. This finding has direct implications for applying CDT-II to new systems: researchers should prioritize genes whose perturbation effects are reproducible across independent experiments. The elimination of RNA and protein language model dependencies—a key design change from CDT v110—enables this flexibility by allowing researchers to work directly with raw per-cell expression rather than fixed embeddings.
Because per-cell RNA expression serves as model input, one might ask whether the model trivially reconstructs its input. This concern reflects a misunderstanding of CDT-II’s task design. Crucially, the input is expression levels (cellular state) while the output is expression changes (perturbation effects)—fundamentally different quantities. The model receives no information about the unperturbed baseline, no pre-computed differences, and no indication of which genes will change. Furthermore, the same cell yields different predictions depending on which DNA locus is queried—the RNA input alone is insufficient without understanding its relationship to the perturbation site. To predict how gene B changes when locus A is perturbed, the model must learn the regulatory relationship between A and B; simply memorizing or reconstructing input expression patterns cannot solve this task.
This task design creates a synergy with CDT-II’s architecture. By providing raw expression as input and predicting expression changes as output—without pre-computing differences—we create a learning problem that requires understanding regulatory relationships. Because CDT-II’s architecture is explicitly structured by the central dogma, these learned relationships are channeled into interpretable attention mechanisms. The attention mechanisms—DNA self-attention, RNA self-attention, and DNA-to-RNA cross-attention—correspond to specific biological relationships: genomic interactions, gene co-regulation, and transcriptional control, respectively. The result is a system where improving prediction accuracy and discovering regulatory biology are the same optimization objective. This alignment explains why CDT-II’s attention maps recover known biology without supervision: the task leaves no shortcut.
Three additional observations support this conclusion. First, gene-level holdout ensures that validation targets—including GFI1B and its entire regulatory network—are entirely absent from training; the model cannot have memorized these relationships. Second, adding noisier genes (9,335 versus 2,361) degraded rather than improved performance, the opposite of what would be expected if the model were simply memorizing input patterns. Third, and most strikingly, the biological coherence of unsupervised attention patterns provides direct evidence of genuine learning: the 6.6-fold enrichment for GFI1B targets ($P=3.5\times 10^{-17}$) means that the model independently discovered a regulatory network that was characterized through years of experimental work—without ever being trained on GFI1B perturbation data.
The smooth learning dynamics observed during training provide additional evidence that CDT-II’s architectural inductive bias aligns with biological reality. Throughout training, validation accuracy tracked training accuracy without divergence (train $r=0.65$, val $r=0.64$), indicating that the model learns generalizable patterns rather than memorizing training examples. This behavior contrasts sharply with conventional deep learning, where capacity-rich models typically overfit unless heavily regularized. We attribute this stability to the correspondence between model structure and biological mechanism: because CDT-II’s information flow mirrors the actual central dogma—DNA encoding regulatory potential, RNA reflecting cellular state, cross-attention modeling transcriptional control—the optimization landscape contains a clear path toward solutions that generalize.
A notable feature of CDT-II is its architectural simplicity. The model uses only standard transformer components—self-attention, cross-attention, and feed-forward layers—without specialized modules, custom loss functions, or complex training procedures. This simplicity is deliberate: by constraining the architecture to mirror the central dogma rather than adding complexity, each attention mechanism retains a clear biological interpretation. In contrast, more complex architectures often sacrifice interpretability for marginal performance gains. CDT-II demonstrates that biological insight can emerge from standard components when they are structured to reflect biological organization. This approach parallels the interpretability-by-design principle demonstrated by BPNet24 for transcription factor binding, extended here to multi-modal regulatory network inference.
Several limitations should be noted. CDT-II’s pseudo-bulk correlation (mean $r=0.84$) surpasses the inter-experimental reproducibility ceiling ($r\approx 0.79$ between Morris et al. and Gasperini et al. screens10,12,14), suggesting that the remaining per-cell prediction gap ($r=0.64$) largely reflects irreducible single-cell stochasticity and inter-experimental variability rather than model limitations. While in silico knockdown experiments—setting gene expression to zero—yielded weak correlations ($r\approx 0.07$), this reflects the challenge of extrapolating far outside the training distribution rather than a failure to learn regulatory relationships. In contrast, gradient-based attribution ($\partial\text{output}_{j}/\partial\text{input}_{i}$), which measures local sensitivity within the training distribution, achieved $r=0.83$ correlation with experimental effect sizes. This demonstrates that CDT-II accurately captures regulatory relationships when queried appropriately, and suggests that gradient-based virtual screening—predicting the direction and relative magnitude of perturbation effects—may be more practical than simulating complete knockdowns. The 2,361-gene set excludes some known GFI1B targets (INSIG1, NPM1, BNIP3), and all experiments were conducted in K562 cells from a single dataset. Finally, attention weights require careful interpretation25,26; we mitigate this concern through external validation (6.6-fold enrichment) rather than relying on attention as post-hoc explanation, but attention should not be interpreted as causal without experimental follow-up. We note that direct comparison with perturbation prediction methods such as GEARS6,27 and CPA7 is not applicable: those methods aim to predict the effects of unseen perturbations without experimental data, whereas CDT-II uses experimental perturbation data to reveal interpretable regulatory structure—fundamentally different objectives requiring different evaluation criteria. Similarly, network inference methods such as SCENIC28 and CellOracle29 reconstruct regulatory networks from observational data; CDT-II complements these by deriving regulatory structure from perturbation responses, grounding network edges in causal experimental data.
A distinctive feature of the AI microscope framework is that its outputs generate experimentally testable hypotheses. The CTCF binding sites and enhancer elements identified by cross-attention (Fig. 6) represent specific genomic loci that can be validated through targeted CRISPRi experiments, and the gene regulatory edges recovered by RNA self-attention (Fig. 5) predict specific regulatory relationships testable by individual gene knockdowns. This creates a feedback cycle: experimental validation of attention-derived hypotheses not only tests the model but also generates new training data that can improve its resolution. As genomic foundation models continue to advance, CDT-II provides a means for experimentalists to interrogate, validate, and refine these models’ representations through their own data—closing the loop between computational prediction and experimental biology.
Looking ahead, CDT-II opens several avenues for biological discovery. First, gradient-based virtual screening ($r=0.83$ with experimental effects) could enable computational prioritization of perturbation experiments: by computing input gradients for any genomic locus in a trained model, researchers could rank regulatory targets by predicted effect magnitude before committing experimental resources. Second, the growing availability of genome-scale perturbation atlases30 across multiple cell types provides an immediate path for testing whether regulatory networks are cell-type-specific or conserved, potentially revealing shared regulatory modules across lineages. Third, integration of additional modalities—scATAC-seq for chromatin accessibility, DNA methylation for epigenetic state, or spatial transcriptomics for tissue context—could further sharpen the microscope’s resolution by providing the model with richer cellular context. Fourth, CDT-II’s modular architecture allows the DNA embedding component to be upgraded as genomic foundation models improve; Enformer could be replaced with more recent models such as AlphaGenome21 or Evo22, potentially improving the resolution of DNA-mediated regulatory predictions without retraining the entire system. Finally, because CDT-II learns from perturbation data rather than observational correlations, it may capture causal regulatory relationships that could inform therapeutic target identification in disease contexts.
CDT-II reframes the central question of biological AI from “what does the model predict?” to “what does the model reveal?” By grounding the architecture in the central dogma and using prediction as calibration rather than objective, CDT-II establishes a paradigm in which AI systems serve as observational instruments for biology—complementing, rather than replacing, the expertise of the researcher. The microscope did not replace the biologist’s eye; it extended what the eye could see. CDT-II aims to do the same for regulatory networks that are invisible to any single experimental assay.
Online Methods
Data sources and preprocessing
Primary CRISPRi dataset. We used the STING-seq v2 dataset12 (GEO accession GSE171452), which profiled 60,505 K562 cells using single-cell RNA sequencing combined with CRISPR interference11,31. Perturbations targeted 447 genomic loci: 27 transcription start sites (TSS) and 420 single-nucleotide polymorphism (SNP) loci. Because the screen is highly multiplexed (mean 93 guides per cell), most cells lacked sufficient guide RNA signal for unambiguous assignment; cells were retained only if exactly one target gene had UMI $\geq 50$ with no competing signal, yielding 10,328 assignable cells (8,250 TSS-perturbed and 2,078 NTC) from the original 60,505. Non-targeting control (NTC) cells (n=2,078n=2{,}078) served as the unperturbed baseline. An additional 7,407 SNP-perturbed cells (420 loci) were used with guide assignments from the original publication12.
Gene set curation. The 2,361-gene set was derived by intersecting genes detected as expressed in both an independent K562 CRISPRi screen14 and the primary STING-seq dataset12. This cross-dataset reproducibility filter yielded 2,360 genes. GFI1B was added as a 2,361st gene based on its established role as a master transcriptional regulator with hundreds of characterized trans-regulatory targets across the transcriptome in the primary dataset12, providing a critical test case for unsupervised network recovery.
Cell-level effect computation. For each perturbed cell $c$ and gene $g$, the log2 fold change was computed as:
$y_{cg}=\log_{2}\frac{x_{cg}+1}{\bar{x}_{g}^{\text{NTC}}+1}$ (1)
where $x_{cg}$ is the CPM-normalized expression of gene $g$ in cell $c$, and $\bar{x}_{g}^{\text{NTC}}$ is the mean CPM expression of gene $g$ across all 2,078 non-targeting control cells. This per-cell computation preserves individual cell-state variation rather than collapsing to pseudo-bulk effect sizes.
DNA embeddings. Pre-computed Enformer3 embeddings of shape [896, 3,072] were generated for each perturbation locus, centered on either the TSS or SNP position, spanning approximately 115 kilobases at 128-base-pair resolution.
Data splits. Training comprised 13,620 cells: 6,213 TSS-perturbed cells (22 genes) and 7,407 SNP-perturbed cells (420 loci). Validation comprised 2,037 cells from five TSS-perturbed genes held out at the gene level: GFI1B (477 cells), TFRC (847 cells), TNFSF9 (349 cells), CD44 (233 cells), and CD52 (131 cells). Gene-level holdout ensures that all cells associated with a validation gene are excluded from training, preventing information leakage.
Model architecture
The CDT-II architecture (Fig. 1) follows the design principles established in CDT v110, with key modifications for raw expression input. The RawExpressionEncoder combines learned gene identity embeddings with projected log1p(CPM) expression values to produce gene-level representations. Enformer DNA embeddings are projected from 3,072 to 512 dimensions. All attention operations use standard scaled dot-product attention13 with 8 heads, embedding dimension $d=512$, FFN hidden dimension 4d=2,0484d=2{,}048, GELU activation, and dropout $p=0.3$. DNA representations pass through 2 self-attention layers; RNA representations pass through 1 self-attention layer followed by 1 cross-attention layer (RNA queries, DNA keys/values). The Virtual Cell Embedder10 uses 4-head attention pooling to compress each modality into a fixed-dimensional vector, which are concatenated and fused via MLP. A two-layer task head predicts log2 fold changes for all 2,361 genes. Total parameters: $\sim$21 million.
Training procedure
CDT-II was trained to minimize mean squared error (MSE) between predicted and observed per-cell log2 fold changes:
$\mathcal{L}=\frac{1}{|\mathcal{B}|}\sum_{c\in\mathcal{B}}\frac{1}{G}\sum_{g=1}^{G}(\hat{y}_{cg}-y_{cg})^{2}$ (2)
where $\hat{y}_{cg}$ is the predicted log2 fold change for gene $g$ in cell $c$, $y_{cg}$ is the observed value, G=2,361G=2{,}361, and $\mathcal{B}$ is a mini-batch. We used the AdamW32 optimizer (learning rate $1\times 10^{-4}$, weight decay $1\times 10^{-5}$) with ReduceLROnPlateau scheduling (factor 0.5, patience 10 epochs, monitoring validation loss). Gradient norms were clipped to a maximum of 1.0. Batch size was 64. Training was performed on a single NVIDIA A100 GPU (40 GB).
Attention map extraction
RNA self-attention. Attention weights $n_{\text{genes}}\times n_{\text{genes}}$] were extracted from the RNA self-attention layer for each cell. For genome-wide analyses, attention weights were averaged across all 8 heads and across validation cells. For cell-to-cell variation analyses, per-cell attention vectors for specific query genes (e.g., GFI1B) were compared using Pearson correlation.
DNA-to-RNA cross-attention. Attention weights $n_{\text{genes}}\times 896$] were extracted from the cross-attention layer, representing each gene’s attention to the 896 genomic positions. Gene-gene similarity was computed as the cosine similarity between pairs of cross-attention vectors, yielding an indirect gene-gene regulatory similarity matrix based on shared genomic attention patterns.
ENCODE enrichment analysis
Cross-attention profiles were compared against ENCODE K562 regulatory annotations20 to assess whether attention autonomously focuses on known regulatory elements. Five ENCODE K562 peak-call datasets were used: CTCF ChIP-seq (ENCFF769AUF), H3K27ac ChIP-seq (ENCFF864OSZ), H3K4me1 ChIP-seq (ENCFF135ZLM), H3K4me3 ChIP-seq (ENCFF313FYW), and DNase-seq (ENCFF422QRZ).
Enrichment testing. For each held-out gene, the mean cross-attention profile (averaged across validation cells and 8 attention heads) was computed over the 896 genomic bins. Bins in the top 10% of attention weight were designated as “high-attention.” Each ENCODE peak file was intersected with the 896-bin coordinate system, and Fisher’s exact tests were computed for each gene–mark combination (25 tests: 5 genes $\times$ 5 marks). Odds ratios were Haldane-corrected (adding 0.5 to each cell of the $2\times 2$ table) to handle zero cells and provide finite 95% confidence intervals. All 23 significant $P$ values were below $1\times 10^{-3}$, well within the Bonferroni-corrected threshold of $0.05/25=0.002$.
Circular permutation test. To control for spatial autocorrelation in both attention profiles and ENCODE peaks, we performed circular permutation tests (n=1,000n=1{,}000). In each permutation, the attention profile was circularly shifted by a random offset along the genomic axis, preserving the autocorrelation structure of the attention signal while disrupting its alignment with ENCODE peaks. The observed enrichment was compared against the null distribution to obtain permutation $P$ values.
Bin classification and effect sizes. Each genomic bin was classified into one of four categories based on ENCODE annotations: promoter (H3K4me3+), active enhancer (H3K27ac+ & H3K4me3-), CTCF-only (CTCF+ & H3K4me3- & H3K27ac-), or unannotated (no peaks). Attention weights were compared across bin classes using Kruskal–Wallis tests. Pairwise effect sizes between each annotated class and unannotated bins were quantified using Cohen’s $d$.
Network analysis
Graph construction. Gene-gene networks were constructed from attention matrices by retaining edges above the top 5% attention weight threshold.
Community detection. Louvain community detection16 (resolution parameter = 1.0, the default) was applied to identify modules of co-regulated genes. For RNA self-attention, this yielded 5 communities; for cross-attention similarity, 53 communities (reflecting finer-grained structure).
GO enrichment. Gene Ontology17 and pathway enrichment was performed using the Enrichr33 API against Gene Ontology Molecular Function, KEGG19, and Reactome18 databases. $P$ values were corrected for multiple testing using the Benjamini–Hochberg procedure.
Hypergeometric test. Overlap between top $N$ attention-ranked genes and top $N$ experimentally ranked genes (by absolute mean log2FC upon GFI1B perturbation) was evaluated using the hypergeometric test: $P(X\geq k)$ where M=2,361M=2{,}361 (universe), $K=N$ (top experimental targets), $n=N$ (top attention targets), and $k$ = observed overlap. The enrichment fold change was computed as $k/(N^{2}/M)$. This symmetric comparison avoids the need to define an arbitrary significance threshold for “known targets.”
Gradient analysis
Input gradients $\partial(\text{output}_{j})/\partial(\text{input}_{i})$ were computed for each gene pair using backpropagation through the full model. Gradient magnitudes were averaged across validation cells to produce a gene-gene gradient matrix complementary to the attention-derived network. To validate that gradients capture regulatory relationships, we compared gradient-derived rankings with experimental effect sizes (mean $|\log_{2}\text{FC}|$ per gene pair). This yielded Pearson $r=0.83$, indicating that CDT-II’s gradients accurately predict the magnitude of regulatory effects. Comparison between gradient and attention rankings used Pearson correlation and top-$N$ overlap analysis, revealing complementary but correlated views of regulatory structure ($r=0.27$ between methods).
Ablation study
Two principal training configurations were compared. The full gene set configuration used 9,335 genes detected in at least 10% of cells in the primary dataset12 (excluding mitochondrial, ribosomal, and hemoglobin genes), applying a single-dataset expression threshold. The filtered gene set configuration used 2,361 genes derived from a cross-dataset reproducibility filter (intersection of two independent CRISPRi screens12,14) plus GFI1B. Both configurations used identical cell splits (13,620 training, 2,037 validation), architecture template, and hyperparameters; parameter counts differed due to gene set size ($\sim$21 million vs $\sim$54 million). Learning curves were compared on the same validation cell set using Pearson $r$.
Code and data availability
CDT-II source code is available at https://github.com/nobusama/CDT2. Enformer embeddings, processed training data, and trained model weights are available at https://huggingface.co/datasets/nobusama17/CDT2-data. The STING-seq v2 dataset is available from the original publication12. CDT v1 is described in ref. 10.
Acknowledgements
The author thanks the developers of Enformer, the Morris laboratory for making the STING-seq dataset publicly available, and the open-source communities behind PyTorch and related tools. Claude (Anthropic) was used for manuscript proofreading and editing. The author assumes full responsibility for the accuracy and integrity of all content in this manuscript.
Funding
This work received no external funding.
Author contributions
N.O. conceived the study, developed the model, performed all analyses, and wrote the manuscript.
Competing interests
The author declares no competing interests.
Supplementary Information
Figure 7: Supplementary Figure 1: Statistical robustness of ENCODE enrichment. (a) Circular permutation test (n=1,000n=1{,}000): observed enrichments (red lines) versus null distributions for GFI1B across three marks (DNase, CTCF, H3K4me3), confirming robustness against spatial autocorrelation artifacts. (b) Threshold robustness: enrichment odds ratios remain significant across attention thresholds from top 5% to top 20%, indicating that results are not dependent on a single arbitrary cutoff.